OpenStudy (anonymous):

x+2y=16 -x+2y=8 solve by elimination method

OpenStudy (anonymous):

y=6

OpenStudy (anonymous):

x+2y=16 -x+2y=8 the x's cancel out you get 2y=16 2y=8 Add both equations 4y=24 then divide both sides by 4 y=6

OpenStudy (anonymous):

ok then i would plug 6 into the x equation right?

OpenStudy (anonymous):

-x+2(6)=8

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

-x+12=8 x =-4?

OpenStudy (anonymous):

no for -x+2y=8 when solve for x (y=6) we get x=4

OpenStudy (anonymous):

-x+2y=8 -x+2(6)=8 -x+12=8 subtract 12 from both sides -x=-4 then divide by -1 both sides. since negative x x=4

OpenStudy (anonymous):

and second equation x+2y=16 x+2(6)=16 multiply 2*6=12 x+12=16 subtract 12 from both sides x=4

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

x+2y=16 -(i -x+2y=8 -(ii x+2y=16 -x+2y=8 [+x and -x cancelled] then , 4y=24 [adding 2y and 2y ; and 16+8] or y=24/4 therefore, y=6 Again keeping value of y in equation i, we get, x+2y=16 or x+2*6=16 or x +12=16 or x=16-12 therefore, x=4